4. You roll a standard die 24 times. Knowing that the event satisfies the requirements for a point) binomial distribution, find the probability that exactly 4 of the rolls show a 6. O 0.214 0 1.000 0.167 0.138 5. You flip a coin 10 times. Knowing that the event satisfies the requirements for a binomial point) distribution, find the probability that exactly 7 of the outcomes are heads 00.120 0.001 0.500 6. A survey is given to 50 people to determine if they prefer white or brown rice at a local ( point) Chinese restaurant. Given the following reasons, determine which fulfills one of the requirements for a binomial distribution. The study has a finite number of trials. O The study is conducted randomly O The survey focuses on food only from one restaurant. O The study is conducted locally
Answer #1

Ans:

4)

Probability of showing a six=1/6

n=24,p=1/6

P(x=4)=24C4*(1/6)4*(5/6)20=0.214

5)Binomial distribution with n=10 and p=0.5

P(7 heads)=10C7*0.57*0.53=0.117

6)

The study has a finite number of trials.